The length of a rectangle flower garden is three feet more than its width. The area is 70 square feet. Using an algbriac equation find the length and the width of the flower garden.
Would I use width = x and length = x+3 = to 70? Is that how I would set this equation up?

Answer by nerdybill(7384)

The length of a rectangle flower garden is three feet more than its width. The area is 70 square feet. Using an algbriac equation find the length and the width of the flower garden.
.
Let x = width
then because "length of a rectangle flower garden is three feet more than its width" we get
x+3 = length
(basically, what you had)
.
Now, you have to apply your knowledge of the area of any rectangle:
area = width * length
giving you your equation:
70 = x(x+3)
now we solve for x:
70 = x(x+3)
70 = x^2+3x
0 = x^2+3x-70
0 = (x+10)(x-7)
x = {-10, 7}
we can throw out the negative solution (extraneous) since it doesn't make sense leaving:
x = 7 feet (width)
.
Length:
x+3 = 7+3 = 10 feet